OpenStudy (anonymous):

2x+6 -6x =30 a.-6 b.3 c. -3 d. 6

OpenStudy (gorv):

-4x=24 x=-6

OpenStudy (anonymous):

8x+6=30 8x=24 x=3

OpenStudy (gorv):

@akc1999

OpenStudy (anonymous):

-6=x

OpenStudy (solomonzelman):

@hartnn, is this right?

OpenStudy (solomonzelman):

Why are they giving work and answers?

OpenStudy (anonymous):

I added the 6 instead of subtracting it but you can do that right?

OpenStudy (pixiedust1):

Okay, let's combine terms. 2+-6 = -4

OpenStudy (pixiedust1):

Now we have -4x+6=30

OpenStudy (pixiedust1):

-4*-6=24 24+6=30

OpenStudy (pixiedust1):

Therefore, the answer is -6

OpenStudy (pixiedust1):

x=-6

OpenStudy (pixiedust1):

Do you understand?